327 band RFI measurements.

last updated 16jan19

dynamic spectra by month for p1693 pulsar search data.

16jan19: 327 Mhz hi freq resolution az swings taken 10jan19.
10jan19: dynamic spectra by month for p1693 pulsar search data.
01jun17: rfi from sband console in dome.
11oct16: az encoder rfi measured in screen room.
04feb16: turning the sband console on,off after removing external ups, power supply
03feb16: 4Mhz comb from sband console coming from ups outside of cabinet
29jan16: cleaning up the 327mhz band.. jan16
28jan16: turning the sband console (upstairs) on,off after the cabinet repairs
28jan16: turning the sband console (upstairs) on and off.. before repairs.
28jan16: fixing the shielding gasket on the turret control box (in the dome)
27jan16: probing the sband console (upstairs) with the anritsu
25jan16: hi frequency resolution azimuth swings
12jun15: anritsu spectrum analyzer in the if/lo rack in the dome.
04mar14: 327 rfi from meteor camera on antenna test shack.
27feb14: 320-325 Mhz drifting birdie seen feb14
23feb14: rfi seen in the p1693 pulsar search
11jul13: P2717 327 incoherent search with puppi show avg spectra, rms, and rfi.
04jun13:measure 3rd harmonic of 102.9,104.1 with a dipole.
30may13: 3rd harmonics and intermods of fm 102.9,104.1 from route 10
24mar13:Mapping screen room birdie strength to detectability on the telescope
22mar23: elo touch screen cash register. Emissions in screen room and on telescope.
28jan13: spectrum analyzer at building 2 a/c unit shows rfi.
24jan13: 327 rfi when building 2 power cycled.
30apr12: 3rd harmonics of fm stations 104.1,102.9 from route 10.
03nov11: possible 2 Mhz comb in 327 Mhz band.
03nov11: check alfa motor controller rfi after cable shielding.
01nov11: power shutdown of visitors center.
10aug11: short duration broadband rfi seen by p2558 in the 327 rcvr.
10aug11: chris salters list of rfi at 327,430 using telescope data.
10jun11: tested compressor monitor on telescope 06jun11 using 327
09mar11: az dependence of 327 rfi
02mar11: alfa motor, cameras, and distomat rfi.
09aug09: rfi from azimuth encoders
11jun09: rfi from sband  console/alfa motor.
18oct06: looking at rfi created by axis 212 web camera.
03oct06: 614Khz comb (encoder combs)
18apr06: the 750 Khz and 40 Khz 327 combs come from the sband console in the dome.
16apr06: 750 Khz and 39.45 Khz combs in 327 receiver
13apr06: testing the trenddata gigabit switch for rfi.
28mar06: look at new ac units in dome. Compare with comb from old unit.
16jan06: 28oct10 az swings show birdies with 1 min periodicity
15jan06: wide band look at Dome ac birdies in 327 Mhz rcvr.
14jan06: 3rd harmonics of fm stations appearing in band
10jan06: placing hipass, bandpass filters before the 1st amp
02nov05 cleaning up the 327 Mhz band.
28oct05: 327 Mhz birdies and their azimuth dependence.
21oct05: 327 Mhz spectral density at output of 1st amplifier (before the rf filter)
16oct05 1 Mhz comb in 327 receiver from both dome AC units
03jun05 16Khz comb in 327 receiver from alfa motor.
03dec03 327 Mhz comb from sband Tx Rack in dome
24nov03 comb at 327 Mhz.
24jul02 2 MHz comb in 327 Mhz receiver
01dec03 327 Mhz comb coming from sb Tx room in dome.

04mar14: meteor camera rfi

    The meteor camera atop the old antenna test shack was found to make rfi at 324 Mhz.
We turned the camera off around 16:07 04mar14.
p1693 was taking 327 drift data when we did this:
The dynamic spectra shows the data during the turn off (.gif)
You can also see this in the 04mar14 dynamic spectra for p1693. (in this image it occurs at 16.75)

processing: x101/140304/metcamrfi.pro

03nov11: possible 2Mhz comb in 327  receiver.

    A 110 second median spectra was taken with the 327 Mhz receiver. The setup and processing was:
The plots show the spectra for the 110 second integration (.ps) (.pdf):
processing: x101/111103/alfamot.pro

03nov11: alfamotor (top)

    The alfa motor cables had extra shielding added when alfa was in the lab during  sep11.  on 03nov11 an rfi test was done with the 327 rcvr to check the rfi being emitted by the alfa motor controller. The setup was:
The dynamic spectra show the spectral density vs time:


processing: x101/111103/alfamot.pro

02mar11: alfamotor,cameras, distomat rfi  (top)

    On 02mar11 test were made to check the rfi from the alfa motor controller as well as the dome cameras.
The setup was:
The dynamic spectra shows the rfi vs time (.gif)
The plots show the average spectra for the on and offs (.ps) (.pdf):


processing: 110302/327rfi.pro

11jun09: rfi from sband console, alfa motor (top)

    A strong comb was seen in the 327 rcvr around the end of may09. It had a 750 Khz spacing which pointed to the sband console. Data was taken on 01jun09 while various equipment was turned on/off. The setup was:
    The sequence of equipment on off was (by record number):
The dynamic spectra for the sequence is shown in the img (.gif) :
The next plot shows a spectral average of the sband console on and off (.ps) (.pdf) :

11jun09: after shielding the console:

    Victor and rey worked on shielding the sband console after this data was taken. A second test was done on 11jun09 to see the results. The setup was:
The image shows the 327 dynamic spectra while the sb console was turned on/off (.gif):


processing: x101/090601/327rfi.pro, 090611/chksbrfi.pro

13apr06: testing the trenddata gigabit switch for rfi.  (top)

    A trend data gigabit Ethernet switch (teg-s41sx) was tested for rfi. This switch has 1 fiber port and 4 copper ports. During the testing there was nothing connected to the switch (except the power). In the screened room rfi was seen from the switch beyond 2 Ghz. The strongest birdie was at 476 Mhz. It was stronger than 50 db above the noise floor of thesbt250 spectrum analyzer (the resolution bandwidth was about 300 hz).
    On 13apr06 data was taken with the 327 receiver on the telescope and the interim correlator. The switch was located in the parking lot next to the control room inside the rfi van. It had a clear shot to the telescope through the front window or the van. The azimuth was positioned to az=242, za=18 (the 1400 Mhz birdie from the wapps is a maximum here). The correlator was set to 390 Khz bandwidth and 2048 channels. After hanning smoothing the resolution bandwidth was 380 hz. 180  1 second integrations were made.
    The data taking started with the switch on. It was turned off at second 80. It was turned back on at second 125. It was shut off again at second 166.


processing: x101/060413/rfi.pro

21oct05: 327 Mhz spectral density at output of 1st amplifier (before the rf filter). (top)

    The output of the 327 Mhz receiver after the first amplifier was measured with the Tektronix ybt250 portable spectrum analyzer. The signal  was taken after the first amp and before the rf filter (the filter limits the band to 30 Mhz about 325 Mhz). The telescope was at an az of 270 and a za of 8.46 degrees. Data was taken covering 0 to 1000 Mhz and then 0 to 500 Mhz. At each band a regular sweep and a peak hold of about 1 minute were recorded. This was done for polA and polB. The plots show that television stations are the largest signals in the receiver band (.ps) (.pdf):      On 20oct05 we saw intermittent intermods/harmonics in the 327 Mhz band that included the fm station 107.3  (3rd harmonic at 322) and intermods with a tv station and 107.3 (we heard the fm station in the IF signal).  The power levels measured above do not look like they should be driving the first amp into compression.
    The 1 minute peak hold has values that are up to 10 db stronger than the single sweep values so the levels are time variable. There is most likely an azimuth dependence to the birdies. It still looks like a good idea to put an rf filter before the first amplifier (since the receiver isn't cooled).
processing: x101/051021/327.pro

03dec03 327 Mhz comb from sband Tx Rack in dome.  (top)

        The 327 Mhz comb (with 372 kHz spacing) is coming from the sband transmitter cabinet in the dome. Rey velez noted that the fingers for making contact between the door and the cabinet were not working. Data was taken with the correlator while the sband rack was on and off. 1 second dumps were done using a 24 kHz resolution and a 190 Hz resolution (after hanning smoothing). A bandpass correction was fit to an average of the band passes when the rack was powered off. The plots show the 327 MHz comb when the sband rack was power on and off. The door from the sband transmitter room to the stairwell was open during these test.
processing: x101/031203/doit.pro

24nov03 comb at 327 Mhz.  (top)

    (note: this comb is coming from the sb tx console in the dome)
A comb appeared at 327 Mhz sometime in oct or nov03. Data was taken on 24nov03 with the interim correlator using a 25 Mhz bw, 2048 channels, 1 second sampling, and centered at 327 MHz. An az swing was done from az=290 (za=10) clockwise to an az at 660 using a rate of .4 deg/sec. The az was then swung counter clockwise from 660 to about 540 (where is stopped because the cable car was in use). The plots show the comb at 327 Mhz: The comb spacing is 372 kHz with the strongest values at  745 Khz (actually 371.97Khz generates the comb at 327 MHz). There is variation in the power for each element vs azimuth but it does not look like something that would be coming from a tower far away. The dropouts and increases do not repeat with azimuth implying that they are a time variation rather than an azimuth dependence. The birdie is probably on the platform or even in the dome. It would be interesting to take some data with the doors to the turret room open and then closed.

24jul02 2 MHz comb in 327 Mhz receiver (resolved). (top)

Resolution: Conducting gasket in huffman box degrades with time.

    After turning off a number of equipment it was found that the birdie was coming from the power meter that is used to measure the power in the 430 Mhz wave guide. It is located on the ceiling of the turret room inside a huffman box. It has an rf cable going to it from the wave guide, a fiber optic cable sending the signal downstairs, and an ac power code that goes through an ac power filter at the input.  Inside the box is a pc card with a VtoF converter on it. This card uses a 1 Mhz crystal. There is another of these boxes in the carriage house measuring the  430 Mhz output power. At first the ac power filter was suspected as the culprit. The box in the carriage house was brought down and tested. The main problem was the conducting gasket around the inside top of the huffman box. It was supposed to improve the contact between the top and the box when the top was closed. The gasket was a rubber (neoprene?) material with a conductive coating. Where the gasket was in contact with the box (top or bottom) the gasket was still conducting. The edge of the gasket had a dark strip which was no longer conducting so the top of the gasket was no longer electrically in contact with the bottom. When the box was closed the gasket was acting as an insulator for the top. The guess is that the part of the gasket in contact with the air oxidized the conducting material. This is a serious problem since most of the huffman boxes we have use this gasket material. It is probably why the color camera birdie was not there when the cameras were installed in their new shielded boxes and then reappeared 6 months later.

01dec03 327 Mhz comb coming from sb Tx room in dome.

    On 01dec03 data was taken with the 327 receiver while tracking blank sky. The correlator was set for 2048 channels over 25 MHz (centered at 327 MHz) and 2048 channels over 195 kHz (centered at 321.7 Mhz. The data was dumped at 1 second intervals and hanning smoothed to give resolutions of 24 kHz and 190 Hz. While the data was being taken, the doors to the sband transmitter room was opened and closed as well as the door to the turret room. Data was taken for 2200 seconds.
    We knew that the signal was coming in through the horn (since it went away when put on load). The doors should give some shielding to birdies generated inside the rooms. If the birdie strength increased when the door was opened, then it was probably coming from inside the room. The plots shows the 327 comb data while the room doors where opened and closed.
processing: x101/031201/doit.pro

rfi measurements home.